drawn from origin are two mutually perpendicular straight lines forming an isosceles triangle together with th straight line 2x+y=a. then the area of the triangle is in terms of 'a'? 


(A)a2/2


(B)a2/3


(C)a2/5


(D)none


 

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear tejas

length of perpendicular on the line 2x+y=a from the origin is


L = |a|/√5

we know that in right angle triangle mid point of hypotnous is  equidistance from all the vertex.

so length of the base = 2L

so area  = 1/2 *L *2L

             =L2 = a2/5

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Arnav jain
7 years ago
Can you please tell the rough figure for this question with proper labelling of sides as i am not able to get it